under a pressure of 14nm-2, some air has a volume of 1.5m3. determine its volume when its pressure is 10nm-2. assuming the temperature is kept constant
1
Expert's answer
2020-06-05T11:53:05-0400

Given data

Initial pressure is "P_1=14 N\/m^2"

Final pressure is "P_2=10 N\/m^2"

Initial volume is "V_1=1.5 m^3"

According to Boyle's law,

"P_1V_1=P_2V_2"

The final volume is

"V_2=\frac{P_1V_1}{P_2}=\frac{(14 N/m^2)(1.5 m^3)}{10 N/m^2}=2.1 m^3"